OpenStudy (anonymous):

what is the geometric mean of 7 and 12??

OpenStudy (mathstudent55):

The geometric mean of a and b is x in this equation: \(\dfrac{a}{x} = \dfrac{x}{b} \) To solve for x, cross multiply and take the square root: \(x^2 = ab\) \(x = \sqrt{ab} \) In your case, a = 7 and b = 12.

OpenStudy (anonymous):

but the main part that gets me confused is the actual square root part.. i have a multiple choice answer... A) 1 5/7 B) 9.5 C) 2 squareroot of 2 D) 4 squareroot of 21 @mathstudent55

OpenStudy (anonymous):



OpenStudy (anonymous):

so how do you simplify the square root of 84? @hl147

OpenStudy (anonymous):

first do any perfect squares fit into 84? for example 4,9,16,25,etc

OpenStudy (anonymous):

hint** 4 fits into 84, 21 times